# LEMMA: Learning from Errors for MatheMatical Advancement in LLMs

### v1

@@ 5.3 Ablation study
We conduct ablation studies to assess the contributions of each component within LEMMA. During the erroneous step collection phase, we exclude the error augmentation module, relying solely on errors generated by the model itself, which we denote as “w/o Error Aug.” To ensure that any decline in performance is not merely due to a reduced sample size, we generate the same amount of revision trajectories as LEMMA, which we refer to as “w/o Error Aug. (90k).” We also perform ablation on the error correction strategy by removing the “Fresh & Restart” method from the revision process, labeled as “w/o Fresh & Restart” and “w/o Fresh & Restart (90k)”. We report the accuracy on in-distribution tasks in Tab.4; for out-of-distribution performance, please refer to Tab.9 in Appendix A.1. It is evident that removing the “error augmentation module” results in a significant performance drop. This decline is not due to sample size, as “w/o Error Aug (90k)” also exhibits a 6.2% accuracy decrease in performance on MATH compared to LEMMA. We attribute this decline to the reduced diversity of error steps, as the model relies solely on errors generated by the student model itself. In contrast, the error augmentation module introduces a variety of meaningful errors, enhancing the model’s ability for reflection and self-correction. Furthermore, excluding the “Fresh & Restart” strategy degrades performance. This decline highlights the essential role of the “Fresh & Restart” correction: by enabling the model to reset and reassess problem-solving pathways, it significantly enhances mathematical reasoning capabilities.
@@ 6 Conclusion
In this work, we introduce LEMMA, a novel framework designed to enhance the mathematical reasoning capabilities of LLMs by systematically learning from errors. Based on a comprehensive analysis of error types, LEMMA employs an error-type grounded mistake augmentation strategy and constructs diverse revision pathways using both the Fix & Continue and Fresh & Restart correction strategies. This framework allows models to autonomously detect and correct errors during the generation process, thereby improving their mathematical reasoning abilities. Extensive experiments demonstrate that LEMMA significantly outperforms SOTA baselines.
@@
While LEMMA represents a significant advancement in enhancing the mathematical reasoning capabilities of large language models, several limitations persist. Firstly, its focus has been solely on mathematical reasoning tasks, leaving its effectiveness and adaptability in other domains unexplored. Moreover, the synthesized dataset used in LEMMA comprises fewer than $90,000$ examples, which is relatively small compared to data augmentation methods like MetaMath. This raises questions about whether an increase in dataset size could continue enhancing the performance. Investigating the synthesis of additional data to push the performance boundaries of LEMMA could be a future work.
@@ A.2.2 Full Error Taxonomy
We present the full error taxonomy in Table 10. Building upon the taxonomy proposed by Li et al. (2024e), we introduce additional error categories to enable a more granular identification of error types. Specifically, we add “Question Misinterpretation Error (QM)”, “Confusing Concept Error (CC)”, and “Nonsensical Output (NO)” to better capture the diverse range of errors that can occur during mathematical reasoning. The expanded taxonomy provides a structured framework for systematically categorizing and addressing the various types of errors encountered in mathematical problem-solving.
@@ A.4 Prompt
Table 10: Error taxonomy adapted from Li et al. (2024e). We add “Question Misinterpretation Error (QM)”, “Confusing Concept Error (CC)”, and “Nonsensical Output (NO)” to allow for a more fine-grained identification of error types. Corresponding examples are provided in Figure 12.

### v2

@@ 5.3 Ablation study
w/o Truncation 69.4 22.8
We perform ablation studies to evaluate the contribution of the two core innovations of LEMMA: (1) the error-type grounded mistake augmentation method (Sec.3.3) and (2) the mixed correction strategies (Sec.3.4). To evaluate the effectiveness of the mistake augmentation module, we exclude it and rely solely on errors generated by the model itself, which we denote as “w/o Error Aug”. To ensure that any performance degradation is not simply attributable to a smaller sample size, we generate the same number of revision trajectories as LEMMA, which we refer to as “w/o Error Aug (90k)”. We also perform ablation on the error correction strategy by removing the “Fresh & Restart” method from the revision process, labeled as “w/o Fresh & Restart” and “w/o Fresh & Restart (90k)”. Lastly, “w/o Truncation” means that we do not truncate the flawed trajectory at the first identified erroneous step, which aims to verify the necessity of the early truncation of the erroneous reasoning trajectory. We report the accuracy on in-distribution tasks in Tab.5; for out-of-distribution performance, please refer to Tab.11 in Appendix A.1. It is clear that removing the “error augmentation module” results in a significant performance drop. This decline is not due to sample size, as “w/o Error Aug (90k)” also exhibits a $6.2\%$ accuracy decrease in performance on MATH compared to LEMMA. We attribute this decline to the reduced diversity of error steps, as the model relies solely on errors generated by the student model itself. In contrast, the error augmentation module introduces a variety of meaningful errors, enhancing the model’s ability for reflection and self-correction. Furthermore, excluding the “Fresh & Restart” strategy degrades performance. This decline highlights the essential role of the “Fresh & Restart” correction: by enabling the model to reset and reassess problem-solving pathways, it significantly enhances mathematical reasoning capabilities.
@@ 6 Conclusion
In this work, we introduce LEMMA, a novel framework designed to enhance the mathematical reasoning capabilities of LLMs by systematically learning from errors. Based on a comprehensive analysis of error types, LEMMA employs an error-type grounded mistake augmentation strategy and constructs diverse revision pathways using both the Fix & Continue and Fresh & Restart correction strategies. This framework allows models to autonomously detect and correct errors during the generation process, thereby improving their mathematical reasoning abilities. Extensive experiments demonstrate that LEMMA significantly outperforms strong baselines.
@@
While LEMMA represents an advancement in enhancing the mathematical reasoning capabilities of LLMs, several limitations persist. Firstly, the synthesized dataset used in LEMMA comprises fewer than $90k$ examples, which is relatively small compared to data augmentation methods like MetaMath. This raises questions about whether an increase in data size could further enhance the performance. Moreover, as LEMMA fine-tuned models have exhibited preliminary reflection capabilities—akin to the “aha moment” of DeepSeek-R1 Guo et al. (2025), they may be good starting points for rule-based RL training. Exploring the use of LEMMA with rule-based RL could be a future direction.
Acknowledgments
This research was supported by National Key R&D Program of China (2022ZD0160201).
@@ A.2.2 Full Error Taxonomy
We present the full error taxonomy in Table 12. Building upon the taxonomy proposed by Li et al. (2024d), we introduce additional error categories to enable a more granular identification of error types. Specifically, we add “Question Misinterpretation Error (QM)”, “Confusing Concept Error (CC)”, and “Nonsensical Output (NO)” to better capture the diverse range of errors that can occur during mathematical reasoning. The expanded taxonomy provides a structured framework for systematically categorizing and addressing the various types of errors encountered in mathematical problem-solving.
@@ A.4 Prompt
A.5 Comparison with Additional Baselines
Table 6: Performance comparison with additional baselines on GSM8K, MATH using LLaMA3-8B.
Method GSM8K MATH Avg.
LEMA An et al. (2023) 75.1 32.8 53.9
Mistake tuning Tong et al. (2024a) 68.9 25.5 47.2
LLM2LLM Lee et al. (2024) 76.8 33.1 54.9
LEMMA 79.2 38.3 58.7
To further illustrate the advantages of our method, we conduct additional experiments comparing LEMMA with additional baselines. The results are in Tab.6. Among these baselines, An et al. (2023) bears the most similarity to ours, as it synthesizes mistake-correction data pairs by instructing GPT-4 to rectify incorrect reasoning paths generated by the reasoning model. However, it relies on ensembling errors from multiple LLMs to create diverse inaccurate reasoning paths, which can be prone to generating irrelevant inaccurate reasoning paths that the student model would not typically make. Furthermore, their method neglects different correction strategies, which restricts the model’s ability to reflect and self-correction. Our two core innovations—(1) the error-type grounded mistake augmentation method and (2) mixed correction strategies (“Fix & Continue” and “Fresh & Restart”)—represent novel contributions that have not been explored in previous work. To reimplement An et al. (2023), we use the same prompt as them, but change the student model / teacher model from LLaMA2 / GPT-4 to LLaMA3 / GPT-4o to align with our setting. We generate correction data based on the training set of GSM8K and MATH. Tong et al. (2024a) constructs a dataset called “CoTErrorSet” includeing errors and PaLM2-annotated error causes. They also incorporate fixed prefixes ([CORRECT RATIONALE] and [INCORRECT RATIONALE]) in fine-tuning, which they refer to as “Mistake Tuning”. However, their method does not concatenate the incorrect rationale and correct rationale, but instead adds a specific prefix in front of the rationale, which is different from ours. For Mistake-tuning Tong et al. (2024a), since the “CoTErrorSet” in Mistake-tuning does not include any data from MATH and is generated by PaLM2, we regenerate it based on the training set of MATH and GAM8K using GPT-4o (the same as ours) and then perform the mistake tuning. LLM2LLM Lee et al. (2024) iteratively synthesizes more data based on the data points that the student model fails to answer. Their focus is on sample-level augmentation through the synthesis of additional data points. In contrast, our method implements sequence-level augmentation by enriching each data point with common errors and subsequent reflection. We also reimplement LLM2LLM Lee et al. (2024) using GPT-4o as the teacher model to ensure a fair comparison. The results in Tab.6 further illustrate the advantages of our method over these baselines.
@@ A.5 Comparison with Additional Baselines
Table 12: Error taxonomy adapted from Li et al. (2024d). We add “Question Misinterpretation Error (QM)”, “Confusing Concept Error (CC)”, and “Nonsensical Output (NO)” to allow for a more fine-grained identification of error types. Corresponding examples are provided in Figure 12.